A positive point charge q1 at the origin and a negative charge q2 at x=2m. The resultant electric field is 10.8 i N/C at x=1m and -8 i N/C at x=3m. Find q1 and q2.


1
Expert's answer
2022-02-17T04:34:39-0500

We need to establish the two equations to determine the total electric field at x=1m and x=3m. The electric field is equal to "\vec{E_i}=k\dfrac{q_i}{r_i^2} \hat{r_i}". We know that q1>0 and q2<0, k=9X109 Nm2/C2, and for the first case at x=1 m, r1=r2=1m while "\hat{r_1}=\hat{i} \text{ and } \hat{r_2}=-\hat{i}". On the second case at x=3m we have r'1=3m, r'2=1m, and "\hat{r'_1}=\hat{r'_2}=\hat{i}". Using that information we find the total electric field as ET=E1+E2:


  • for x=1m:


"(9\times10^9 Nm^2/C^2)\Big( \frac{q_1}{(1\,m)^2} -\frac{q_2}{(1\,m)^2} \Big)\hat{_i}=(10.8\,N/C) \hat{_i}"

  • for x=3m:

"(9\times10^9 Nm^2/C^2)\Big( \frac{q_1}{(3\,m)^2} +\frac{q_2}{(1\,m)^2} \Big)\hat{_i}=(-8\,N/C) \hat{_i}"


The system that has to be solved is:

"q_1-q_2=1.2\times10^{-9}\,C
\\ q_1+9q_2=-8\times10^{-9}\,C"


After solving, in conclusion, q1 = 2.8 X 10-10 C and q2= -9.2 X 10-10 C.